A hollow sphere of radius 2R is charged to V volt and another smaller sphere of radius R is charged to V/2 volt.Then the smaller sphere is placed inside the bigger 
sphere without changing the net charge on each sphere.The potential difference between the two spheres would be ?

Grade:12th pass

1 Answers

Ravneet
5 years ago
THEORY: inside the hollow sphere the potential at every point is constant and is equal to the potential at the surface 
                   because electric field is zero inside the hollow sphere 
new potential of small sphere when placed inside bigger sphere = potential due to bigger sphere + existing potential of the smaller sphere
new potential of smaller sphere =  V + V/2 = 3V/2 
new potential of bigger sphere  = existing potential of bigger sphere + potential due to the smaller sphere (  at distance 2R potential due to small sphere will be V/4)
new potental of bigger sphere = V + V/4 = 5V/4 
potental difference=new potential of smaller sphere – new potental of bigger sphere
pootential difference = V/4